A second order reaction is 41.5% complete in 500 sec.?

a) Calculate the rate constant? 

b) What is the value of the half-life? 

c) How long will it take for the reaction to go to 25%, 80% completion?


1
Expert's answer
2021-04-09T04:05:13-0400

second order kinetics


"\frac{1}{[A]} = \frac{1}{[A_0]} + kt"


since it's 41.5% complete

so remaining amount is 58.5 %

[A] = 0.585 [A0]


"\frac{1}{[A]} - \frac{1}{[A_0]} = k×500"


"\frac{1}{0.585[A_0]} - \frac{1}{[A_0]} = k×500"


"\frac{1.70-1}{[A_0]} = k×500"


"k = \frac{0.70}{[A_0]×500}"


"k = \frac{0.0014}{[A_0]}"


"t_{1/2} = \frac{1}{k[A_0]}"


"t_{1/2} = \frac{1}{\frac{0.0014}{[A_0]}×[A_0]}"


"t_{1\/2} = 714.2 sec"


for 25%

"\frac{1}{[A]} - \frac{1}{[A_0]} = kt"


"\frac{1}{0.75[A_0]} - \frac{1}{[A_0]} = \frac{0.0014}{[A_0]}×t"


"t = \frac{0.33}{0.0014} = 235 sec"


for 80%

"\frac{1}{[A]} - \frac{1}{[A_0]} = kt"


"\frac{1}{0.20[A_0]} - \frac{1}{[A_0]} = \frac{0.0014}{[A_0]}×t"


"t = \frac{4}{0.0014} = 2857.14 sec"